Determine the pH of an aqueous solution of 0.05M Ba(OH)2
1
Expert's answer
2020-07-10T05:24:23-0400

Ba(OH)2 will completely dissociate into Ba2+ and 2OH- ions

OH- concentration will be double of concentration of Ba(OH)2

so [OH-] = 2*0.05 M = 0.1M

pOH= -log([OH-])

pOH= -log(0.1)

pOH = 1

pH= 14- pOH

pH = 14-1 = 13